Simplify: -48 ÷ (7 – 9)³ - 2[ 1 – 5 (2 – 6) + 3²]
Answer by stanbon(75887)

Simplify: -48 ÷ (7 – 9)³ - 2[ 1 – 5 (2 – 6) + 3²]
------
= -48 / (-2)^2 - 2[1 -5(-4) + 9]
------
= -48 / 4 - 2[1 +29]
------
= -12 - 2(30)
------
= -72
===========
Cheers,
Stan H.
===========